OpenStudy (anonymous):

Find an equation of the plane tangent to the surface e^(xy) + z^2 – z = 1 at the point (0, -1, 1).

OpenStudy (amistre64):

the tangent plane has a remarkable similarity to the point slope form of a line

OpenStudy (amistre64):

i think of it as the point slop slope form of a plane

OpenStudy (amistre64):

\[z-z_o=Fx(x-x_o)+Fy(y-y_o)\]

OpenStudy (amistre64):

or, another way to view it, is:\[Fx(x-x_o)+Fy(y-y_o)+Fz(z-z_o)=0\]

OpenStudy (amistre64):

F(x,y,z) = e^(xy) + z^2 – z = 1 Fx(x,y,z) = ye^(xy) Fy(x,y,z) = xe^(xy) Fz(x,y,z) = 2z – 1 apply your point to those to define them